Find the following values using the equations and then a financial calculator. Compounding/discounting occurs annually. Do not round intermediate calculations. Round your answers to the nearest cent.

  1. An initial $500 compounded for 1 year at 10%.

    $  

  2. An initial $500 compounded for 2 years at 10%.

    $  

  3. The present value of $500 due in 1 year at a discount rate of 10%.

    $  

  4. The present value of $500 due in 2 years at a discount rate of 10%.
Answer #1

Equations:

1.
=500*(1+10%)^1=550

2.
=500*(1+10%)^2=605

3.
=500/(1+10%)^1=454.545454545455

4.
=500/(1+10%)^2=413.223140495868

Financial calculator
1.
N=1
PV=-500
PMT=0
I/Y=10%
CPT FV=550

2.
N=2
PV=-500
PMT=0
I/Y=10%
CPT FV=605

3.
N=1
FV=-500
PMT=0
I/Y=10%
CPT PV=454.545454545455

4.
N=2
FV=-500
PMT=0
I/Y=10%
CPT PV=413.223140495868